The development budget of a municipality in Ghana includes expenditure on Feeder Roads, Schools and Water supply. The expenditure on roads, schools and water supply

are in the ratio 7:15:2. If the expenditure on roads is GH¢28,000,000.00, Find the expenditure on;

(i) Schools

(ii) Water supply

(iii) What is the total budget for these three projects.

(b)The cost of maintaining the municipality libraries is GH¢900,000.00 and this is met from the expenditure on schools. What percentage correct to three significant figures, of the expenditure on schools is spent on maintaining libraries.


C. After making a down payment of GH¢4000 for an automobile, the municipality paid GH¢400 per month for 36 months with interest charged at 12% per year compounded

monthly on the unpaid balance.

(i) What was the original cost of the car?

(ii) What portion of the Municipality’s total car payments went toward interest

charges?


1
Expert's answer
2021-05-07T10:21:32-0400

Let the expenditure on roads, schools and water supply be 7x, 15x,2x respectively.

Expenditure on roads "=7x=28,000,000"


"x=\frac{28,000,000}{4}=4,000,000"


So the expenditure on

(i)schools"=15x=15\times4,000,000"

"=GH\u00a2 60,000,000"


(ii) water supply"=2x=2\times 4,000,000"

"=GH\u00a2 8,000,000"


(iii)total expenditure

"=28,000,000+60,000,000+8,000,000=GH\u00a2 96,000,000"


(b)

Cost of maintaining libraries"=GH\u00a2900,000"

"900,000=x\%of60,000,000"

"\frac{x}{100}\times 60,000,000=900,000"


"x=\frac{900,000 \times 100}{60,000,000}=1.50\%"


(c)


Given



using the excel PV function



(i) Original cost of the car = 16,043

(ii) Portion paid towards interest=2357


Formulas: